OpenStudy (anonymous):

the difference of two numbers is 18 and their sum is 90. Find the two numbers

OpenStudy (anonymous):

54 36

OpenStudy (anonymous):

x-y=18 x+y=90 You can also solve the above system and get your answer

OpenStudy (anonymous):

x=108 and y=90

OpenStudy (anonymous):

acutally when you first add the equztions, you get : 2x=108----->x=54 Then plug in 54 for x in either equation, ill choose the first one:54-y=18---->-y=-36 then divide both side by -1 and get:y=36 x=54 and y=36

OpenStudy (anonymous):

exaclty solve x-y=18 x+y=90 adding both the equations will cancel out the 'y' term n will give 2x=108 \[\implies x=54\] put this in any equation and you will get y=36

OpenStudy (anonymous):

ohh right

OpenStudy (anonymous):

@princess sorry your answer is wrong :(

OpenStudy (anonymous):

yah i know

OpenStudy (anonymous):

x=54

OpenStudy (anonymous):

y=36